OpenStudy (anonymous):

Find the solution of y=6x+1 for x=5 WILL GIVE MEDAL FOR CORRECT ANSWER! Answer Choices: A.(5,2/3) B.(-29,5) C. (5,36) D.(5,31)

OpenStudy (anonymous):

first off. Eliminate choice B because -29 is not 5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

next plug in 5 into the equation and simplify

OpenStudy (anonymous):

y= 6(5) + 1 y=30 + 1 y= 31

OpenStudy (anonymous):

5,31

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

I got another one